OpenStudy (anonymous):

a farmer decides to enclose a rectangular garden using the side of a barn as one side of the rectanlge. what is the max area the farmer can enclose with 100 feet of fence? what should the dimensionsof the garden be to give this area

OpenStudy (anonymous):

25 feet by 25 feet

OpenStudy (anonymous):

the answer says 50 by?feet?

OpenStudy (anonymous):

It's not possible because when you Calculates Perimeter it should be 100 but if we put 50 it becomes 200 ....25 should be the answer

OpenStudy (anonymous):

Panther it says 100 feet is the total length so 100 should be the perimeter 100 = 2x + 2y

OpenStudy (anonymous):

Panther You got it?

OpenStudy (anonymous):

@Ishaan94 "using the side of a barn as one side of the rectanlge."

OpenStudy (anonymous):

yeah sorry ..i did it wrong ..thanks estudier : )

OpenStudy (anonymous):

x is a side, length is 100-2x, A = x(100-2x) = 100x-2x^2 Diff = 0 for max so 4x = 100, side is 25 So area 25*100-2x = 25*50 =1250 sq ft.